# Multimodal 3D Reasoning Segmentation with Complex Scenes

### v3

@@ Abstract
1 Introduction
With the recent advancements in deep learning, 3D scene understanding has become one key component in various real-world tasks such as embodied AI and autonomous driving. The capabilities of understanding spatial relations among 3D objects in scenes and grasping the intention of human users have become essential for machines for interpreting 3D scenes, interacting with 3D objects in scenes, and achieving various complex and real-world missions while navigating within 3D environments.
@@ 1 Introduction
On the other hand, most existing 3D scene understanding work does not possess reasoning and interpretation abilities for interacting with user textual inputs. For example, several studies introduce foundation models such as large language models (LLMs) to empower 3D scene understanding on captioning [7, 5, 14], question answering [2, 28, 31, 10, 14], visual grounding [18, 58, 11, 47, 4, 14, 21], and referring [34, 42, 17, 51], but they are short of reasoning capabilities for deducing human intentions. Several recent studies attempt to introduce the reasoning ability of LLMs into 3D scene understanding tasks, but most of them focus on segmenting single or single-category objects and cannot handle complex scenes with multiple objects of different categories [4, 13, 16]. As a result, they cannot understand 3D spatial relations among objects and produce fine-grained textual explanations, hampering their applications in various real-world scenarios that often come with multiple objects of different categories.
We propose a multi-object 3D reasoning segmentation task that can produce 3D segmentation masks and textual explanations with rich 3D spatial relations among objects in scenes, given 3D scenes and user questions as inputs. The task consists of two key components. The first is ReasonSeg3D, a large-scale and high-quality benchmark that can evaluate 3D reasoning segmentation with multiple 3D objects and rich spatial relations among them. Different from existing 3D reasoning segmentation benchmarks [13, 16], ReasonSeg3D expands the scope into multi-object space which is well aligned with real-world tasks that often come with multiple objects. In addition, ReasonSeg3D integrates 3D spatial information into question-answer pairs, where the 3D spatial relations in textual answers benefit 3D reasoning segmentation clearly. The second is MORE3D, a simple yet effective technique that enables multi-object reasoning segmentation with textual explanations to users’ questions. MORE3D extracts object-specific point cloud embedding for precise prediction of multi-category objects in complex 3D scenes. It incorporates textual explanations with detailed 3D spatial relations of multiple objects in answers, guiding the model toward a comprehensive understanding of 3D scenes. The generated textual answers support accurate segmentation and comprehensive reasoning for complex 3D scenes. As illustrated in Figure 1, MORE3D demonstrates strong reasoning capability for comprehending the intention behind user’s input questions and complex 3D scenes, producing accurate 3D segmentation and explanatory answers with respect to multiple 3D objects in scenes.
@@
2 Related Work
2.1 Language-Instructed 3D Tasks
@@
Integrating point clouds with natural language processing has widespread applications, drawing increasing interest in language-instructed 3D scene understanding. Existing language-instructed 3D scene understanding methods can be broadly grouped into two categories. The first category focuses on 3D segmentation [30, 46, 3, 9, 48, 19, 27, 34, 42, 17, 51, 32, 37, 43, 41], such as OpenScene [32], Openmask3D [37], 3D-STMN [43], RG-SAN [41], which produce segmentation masks but lack reasoning abilities and textual or conversational output. This limits their ability to comprehend the user’s intention and interact with the user in real-world applications. The second category focuses on tasks such as 3D captioning [7, 5, 14], 3D question answering [2, 28, 31, 20, 10, 14], and visual grounding [18, 58, 11, 47, 4, 14, 21, 54, 55, 56]. They can generate textual outputs like phrases or conversational outputs, but leave the fine-grained segmentation task untouched and have no reasoning ability either. Different from existing methods, the proposed MORE3D predicts textual answers with explanation and accurate segmentation of multiple 3D objects within complex scenes, demonstrating strong reasoning capability in comprehending the intention behind the user’s input questions.
2.2 Reasoning Segmentation
Reasoning Segmentation is first introduced by LISA [24] to generate segmentation masks from complex, implicit textual queries. Specifically, LISA integrates LLaVA [26] with SAM, enhancing segmentation through the vision-language model’s reasoning capabilities. Following LISA, PixelLM [35] improves pixel-level reasoning segmentation using multimodal models with a lightweight decoder and segmentation codebook, LLM-Seg [39] bridges the Segmentation Anything Model and LLMs by selecting mask proposals, and LLaVASeg [49] incorporates query-focused segmentation into large language models where chain-of-thought prompting is adopted to preserve dialogue functions. In addition, VISA [45] extends reasoning segmentation to video by combining multimodal language models with a mask decoder, facilitating complex video segmentation from implicit text queries and world knowledge. In the 3D domain, PARIS3D [22] and Reasoning3D [6] focus on part segmentation with explanations for individual 3D objects, leaving 3D segmentation of complex scenes untouched. Recent studies like SegPoint [13] and Reason3D [16] integrate LLMs’ reasoning ability into 3D segmentation, but they are limited to segmentation for single-category objects and have no textual explanations. In contrast, our approach targets multi-object 3D segmentation and provides textual explanations in the output answer, enhancing the model’s understanding of 3D spatial relations and offering a more practical solution for real-world complex scenes.
2.3 Large Multimodal Models
Inspired by the remarkable learning ability of Large Language Models (LLMs), recent research has expanded into the visual domain and developed a series of Large Multimodal Models (LMMs) [1, 52]. The prevalent approach focuses on aligning visual representations with the linguistic embeddings of LLMs. For example, BLIP-2 [25] and mPLUG-OWL [50] encode image features with a visual encoder, integrating them into the LLMs with text embeddings. LLaVA [26] and MiniGPT4 [57] align image-text features followed by instruction tuning, and they also explore image retrieval for LLMs. Recent studies delve into the integration of multimodal LLMs with vision tasks. For example, VisionLLM [40] provides an interface for vision-centric tasks via instruction tuning though it does not exploit LLMs for complex reasoning. VisionLLM-v2 [44] integrates visual perception, understanding, and generation within a unified framework by using a “super link" to connect the multimodal large model with task-specific decoders. DetGPT [33] introduces multimodal LLMs into open-vocabulary detectors for instruction-based detection tasks. GPT4RoI [53] introduces spatial boxes as inputs, training on region-text pairings. LISA [24] enhances segmentation in multimodal LLMs by introducing a <SEG> token. These existing studies primarily target downstream tasks in the 2D domain. In contrast, our approach extends into the 3D domain to enable multi-object segmentation of 3D point clouds and provides textual explanations to support reasoning in the segmentation process.
Figure 2: Illustration of the prompt template in our dataset generation. (a) One example prompt template in our dataset generation for 3D multi-object reasoning. (b) With a sample input image to GPT-4o and the corresponding ground-truth segmentation on the top, the two boxes below present one generated question-answer pair where the text of different colors highlights different objects.
3 ReasonSeg3D Dataset
Most existing reasoning segmentation datasets are not suitable for studying the proposed 3D multi-object reasoning segmentation task. Specifically, existing reasoning segmentation datasets have two critical limitations. First, the existing 2D reasoning segmentation datasets [24, 35, 39, 45] lack 3D data, making them unsuitable for 3D tasks. Second, the existing 3D reasoning datasets [13, 16] focus on single-category objects as illustrated in Table 1, and they contain only a few hundred scenes which are far fewer than standard 3D segmentation datasets [8, 36]. Additionally, these 3D reasoning datasets lack textual explanations with spatial information, hindering them from training MLLMs for better 3D spatial relation understanding. We bridge this gap by proposing a data generation pipeline as well as a new 3D reasoning segmentation dataset ReasonSeg3D, with more details to be elaborated in the following subsections.
Table 1: Comparison of existing 3D scene-level reasoning segmentation datasets. ReasonSeg3D stands out for its large-scale, high-quality data, supporting segmenting multiple 3D objects across multiple categories and offering explanations with 3D spatial relations.
@@ 3 ReasonSeg3D Dataset
# Scene # QA Pair Single Object Multiple Objects Multiple Categories Explanation
Instruct3D [13] ECCV 2024 280 2,565 × ×
Reason3D [16] arXiv N/A 2,484 × × ×
@@
3.1 Dataset Definition
@@
In the proposed ReasonSeg3D, each point cloud $P$ is paired with multiple $\{x_{que},y_{ans},M\}$ triplets, where $x_{que}$ is a question targeting one or more objects in the image, $y_{ans}$ is the textual answer containing the explanation for the reasoning segmentation, and $M$ represents the 3D segmentation masks corresponding to $y_{ans}$. The question $x_{que}$ is designed to require world knowledge and reasoning ability to accurately identify and segment multiple objects. For example, instead of directly asking “Where are the sofa and table?", ReasonSeg3D formulates the question by “Where would be the most suitable place for reading a book in this layout?". An answer $y_{ans}$ in ReasonSeg3D is constructed to involve multiple objects, and it also includes explanations with 3D spatial relations. For example, the $y_{ans}$ to the above question is formulated by “The corner with the sofa and a small table next to it can serve as a perfect reading nook, providing comfort and a quiet atmosphere.".
3.2 Dataset Generation Pipeline
Several reasoning segmentation datasets [24, 39] employ LLaVA [26] for image captioning and text-only GPT-4 to produce question-answer pairs according to generated captions. However, GPT-4 cannot understand the image content well and its generated question-answer pairs often lack crucial 3D spatial relations that are essential for 3D reasoning segmentation in various real-world tasks.
We design a novel data generation pipeline that introduces GPT-4o which possesses superior visual content understanding capabilities. The pipeline allows the generation of more practical questions and answers by incorporating spatial relations among objects in scenes. Specifically, we employ scene image and ground-truth segmentation to enhance GPT-4o’s 3D spatial understanding capabilities. The prompt template is structured in two parts: the first part contains basic requirements for the GPT-4o and the second specifies detailed requirements for generating questions and answers with a focus on describing 3D spatial relations among objects. With this template, GPT-4o autonomously selects objects to form question-answer pairs that reflect the scene’s content and spatial layout. On top of that, we conduct human verification to ensure the quality of the generated question-answer pairs. Figure 2 shows one example of the designed prompt template and one question-answer pair generated by using the template.
3.3 Dataset Statistics
ReasonSeg3D comprises 1513 scenes and 20,113 data samples in total, with point clouds sourced from ScanNetv2 [8]. Following [8], we divide ReasonSeg3D into training and validation sets, comprising 1201 and 312 scenes, respectively. In the generated dataset, each scene has 13.3 questions on average. The dataset contains 20 different object categories for segmentation. Our dataset focuses on indoor scenes, which are critical for a wide range of real-world embodied applications, such as household robotics and augmented reality, where accurate reasoning and perception in complex, cluttered environments are essential.
@@ 3.3 Dataset Statistics
Figure 3: Overview of our proposed MORE3D: Given an input point cloud, the 3D Encoder first extracts per-point features $F_{p}$ and projects them into sequential features $F_{s}$. The sequential features $F_{s}$, together with the textual input $X_{que}$, are then fed into a multimodal LLM to perform reasoning, obtaining LLM embeddings and then producing textual answers $\hat{Y}_{ans}$ with both detailed explanations and descriptions of 3D spatial relations among multiple objects. Finally, after object-specific embedding extraction, object embeddings for multiple <SEG> tokens $F_{seg}$ and the per-point features $F_{p}$ are passed to the 3D Decoder to produce 3D segmentation masks. The module marked with a snowflake icon is frozen during training, while those marked with a flame icon are trainable.
4 Method
4.1 Task Definition
@@
4.2 Overall Framework
Figure 3 shows the framework of the proposed MORE3D which enables Multi-Object 3D Reasoning segmentation and provides textual Explanations based on the user’s input question. Given an input point cloud $P$, the 3D Encoder first extracts per-point features $F_{p}$ and then projects $F_{p}$ into sequential features $F_{s}$ that can be processed by LLMs. The sequential features $F_{s}$ and input text queries are processed by the LLM, obtaining LLM embeddings and then generating textual answers $\hat{Y}_{ans}$ that include textual explanations $\hat{Y}_{text}$ and multi-object <SEG> tokens $\hat{Y}_{\texttt{<SEG>}}$. The <SEG> tokens $\hat{Y}_{\texttt{<SEG>}}$ indicate the request for 3D segmentation masks, and the corresponding object embeddings $F_{seg}$ are extracted from the LLM embeddings. These object embeddings $F_{seg}$ are further combined with per-point features $F_{p}$ within 3D Decoder by dot product, ensuring that the object-specific information interacts with the overall point cloud features to generate the final 3D segmentation masks $\hat{M}$ and classification predictions.
@@ 4.2 Overall Framework
Different from prior studies on 2D reasoning segmentation [24, 35, 39], MORE3D is tailored for 3D tasks and features a 3D-specific encoder-decoder for capturing complex 3D information. In addition, MORE3D extracts object-specific point cloud embeddings to handle multiple objects in 3D scenes, resolving unique challenges in 3D reasoning segmentation effectively. Further, MORE3D generates textual explanations that incorporate 3D spatial relations among objects, enabling explainable and spatially grounded reasoning of 3D scenes.
4.3 Multi-Object 3D Reasoning Segmentation
We design an object-specific embedding extraction approach to achieve multi-object 3D reasoning segmentation. The embedding extraction approach enables precise segmentation and reasoning for multiple objects of different categories in complex point clouds, integrating the LLM with a 3D point cloud decoder to generate segmentation masks and textual explanations.
The LLM takes point cloud features $F_{s}$ and text queries $X_{que}$ as input, generating textual answers $\hat{Y}_{ans}$, which include textual explanations $\hat{Y}_{text}$ and multi-object <SEG> tokens $\hat{Y}_{\texttt{<SEG>}}$. Specifically, each output textual answer consists of a paragraph of textual explanations accompanied by a set of <SEG> tokens, where the number of <SEG> tokens equals the number of 3D objects indicated in the input text query. In each answer, the 3D object’s name is followed by a <SEG> token. For example, the LLM might generate the following answer: “You can use the spacious sofa <SEG> for seating, positioned near a central table <SEG> for drinks and snacks, while additional chairs <SEG> provide extra seating options for guests." Each <SEG> token indicates a request for a point cloud segmentation mask.
@@ 4.3 Multi-Object 3D Reasoning Segmentation
4.4 Explainability
The LLM-generated textual answers $\hat{Y}_{ans}$ contain explanations of the implicit intention of the user’s questions. In addition, the answer includes the descriptions of 3D spatial relations of multiple objects which provide useful guidance for identifying these objects. Unlike most existing reasoning-based segmentation methods [39, 45, 13, 16] that just output statements like "It is <SEG>" without further explanation, our generated answer integrates detailed explanations about 3D spatial relations to enhance segmentation. The ground truth answers $Y_{ans}$ also incorporate these 3D spatial relations and are used to supervise the generated output textual answers, guiding the model to capture spatial information effectively. The textual answer loss is defined by:
@@ 4.5 Training Objectives
where $M$ denotes the ground truth segmentation mask, $\hat{M}$ denotes the predicted segmentation mask, $BCE$ denotes the binary cross-entropy loss and $DICE$ denotes the Dice loss [29]. We use the cross-entropy loss to supervise the point cloud classification as follows:
@@
5 Experiment
5.1 Experimental Settings
@@
Following prior studies on reasoning segmentation [24, 35, 39], we adopt two evaluation metrics including cumulative IoU (cIoU) and generalized IoU (gIoU). cIoU is computed as the cumulative intersection over cumulative union, while gIoU is the average Intersection-over-Union (IoU) across all samples.

### v4

@@ Abstract
Introduction
With the recent advancements in deep learning, 3D scene understanding has become one key component in various real-world tasks such as embodied AI and autonomous driving. The capabilities of understanding spatial relations among 3D objects in scenes and grasping the intention of human users have become essential for machines to interpret 3D scenes, interact with 3D objects in scenes, and achieve various complex and real-world missions while navigating within 3D environments.
On the other hand, most existing 3D scene understanding work does not possess reasoning and interpretation abilities for interacting with user textual inputs. For example, several studies introduce foundation models such as large language models (LLMs) to empower 3D scene understanding on captioning (Chen et al. 2023b, a; Hong et al. 2023), question answering (Azuma et al. 2022; Ma et al. 2023; Parelli et al. 2023; Guo et al. 2023b; Hong et al. 2023), visual grounding (Huang et al. 2022; Zhu et al. 2023; Guo et al. 2023a; Yang et al. 2024a; Chen et al. 2024a; Hong et al. 2023; Kang et al. 2024), and referring (Qian et al. 2024; Wu et al. 2024b; Huang et al. 2021; Yuan et al. 2021), but they are short of reasoning capabilities for deducing human intentions. Several recent studies attempt to introduce the reasoning ability of LLMs into 3D scene understanding tasks, but most of them focus on segmenting single or single-category objects and cannot handle complex scenes with multiple objects of different categories (Chen et al. 2024a; He et al. 2024; Huang et al. 2025). As a result, they cannot understand 3D spatial relations among objects and produce fine-grained textual explanations, hampering application in real-world scenarios with multiple objects of different categories.
@@ Introduction
We propose a multi-object 3D reasoning segmentation task that can produce 3D segmentation masks and textual explanations with rich 3D spatial relations among objects in scenes, given 3D scenes and user questions as inputs. The task consists of two key components. The first is ReasonSeg3D, a large-scale and high-quality benchmark that can evaluate 3D reasoning segmentation with multiple 3D objects and rich spatial relations among them. Different from existing 3D reasoning segmentation benchmarks (He et al. 2024; Huang et al. 2025), ReasonSeg3D expands the scope into multi-object space which is well aligned with real-world tasks that often come with multiple objects. In addition, ReasonSeg3D integrates 3D spatial information into question-answer pairs, where the 3D spatial relations in textual answers benefit 3D reasoning segmentation clearly. The second is MORE3D, a simple yet effective technique that enables multi-object reasoning segmentation with textual explanations to users’ questions. MORE3D extracts object-specific point cloud embedding for precise prediction of multi-category objects in complex 3D scenes. It incorporates textual explanations with detailed 3D spatial relations of multiple objects in answers, guiding the model toward a comprehensive understanding of 3D scenes. The generated textual answers support accurate segmentation and comprehensive reasoning for complex 3D scenes. As illustrated in Figure 1, MORE3D demonstrates strong reasoning capability for comprehending the intention behind user’s input questions and complex 3D scenes, producing accurate 3D segmentation and explanatory answers with respect to multiple 3D objects in scenes.
@@
Related Work
Language-Instructed 3D Tasks
@@
Integrating point clouds with natural language processing has widespread applications, drawing increasing interest in language-instructed 3D scene understanding. Existing language-instructed 3D scene understanding methods can be broadly grouped into two categories. The first category focuses on 3D segmentation (Nguyen et al. 2024; Yan et al. 2024b; Boudjoghra et al. 2024; Ding et al. 2023; Yang et al. 2024b; Huang et al. 2023; Liu et al. 2023b; Qian et al. 2024; Wu et al. 2024b; Huang et al. 2021; Yuan et al. 2021; Peng et al. 2023; Takmaz et al. 2023; Wu et al. 2024c, a), such as OpenScene (Peng et al. 2023), Openmask3D (Takmaz et al. 2023), 3D-STMN (Wu et al. 2024c), RG-SAN (Wu et al. 2024a), which produce segmentation masks but lack reasoning abilities and textual or conversational output. This limits their ability to comprehend the user’s intention and interact with the user in real-world applications. The second category focuses on tasks such as 3D captioning (Chen et al. 2023b, a; Hong et al. 2023; Chen et al. 2024c), 3D question answering (Azuma et al. 2022; Ma et al. 2023; Parelli et al. 2023; Jiang et al. 2025; Guo et al. 2023b; Hong et al. 2023; Chen et al. 2024c), and visual grounding (Huang et al. 2022; Zhu et al. 2023; Guo et al. 2023a; Yang et al. 2024a; Chen et al. 2024a; Hong et al. 2023; Kang et al. 2024; Zhang, Gong, and Chang 2023; Zhang et al. 2024; Zhu et al. 2024a; Chen et al. 2024c). They can generate textual outputs like phrases or conversational outputs, but leave the fine-grained segmentation task untouched and have no reasoning ability either. Different from existing methods, the proposed MORE3D predicts textual answers with explanation and accurate segmentation of multiple 3D objects within complex scenes, demonstrating strong reasoning capability in comprehending the intention behind the user’s input questions.
@@
# Scene # QA Pair S Object M Objects M Categories Explanation
Instruct3D (He et al. 2024) ECCV 2024 280 2,565 × ×
Reason3D (Huang et al. 2025) 3DV 2025 - 2,484 × × ×
@@
Table 1: Comparison on 3D scene-level reasoning segmentation datasets. ReasonSeg3D stands out with large-scale, high-quality data, supporting segmenting multiple 3D objects across multiple categories and offering explanations with 3D spatial relations. S denotes single, and M denotes multiple.
Reasoning Segmentation
Reasoning Segmentation is first introduced by LISA (Lai et al. 2024) to generate segmentation masks from complex, implicit textual queries. Specifically, LISA integrates LLaVA (Liu et al. 2023a) with SAM, enhancing segmentation through the vision-language model’s reasoning capabilities. Following LISA, PixelLM (Ren et al. 2024) improves pixel-level reasoning segmentation using multimodal models with a lightweight decoder and segmentation codebook, LLM-Seg (Wang and Ke 2024) bridges the Segmentation Anything Model and LLMs by selecting mask proposals, and LLaVASeg (Yang et al. 2024c) incorporates query-focused segmentation into large language models where chain-of-thought prompting is adopted to preserve dialogue functions. In addition, VISA (Yan et al. 2024a) extends reasoning segmentation to video by combining multimodal language models with a mask decoder, facilitating complex video segmentation from implicit text queries and world knowledge. In the 3D domain, PARIS3D (Kareem, Lahoud, and Cholakkal 2024) and Reasoning3D (Chen et al. 2024b) focus on part segmentation with explanations for individual 3D objects, leaving 3D segmentation of complex scenes untouched. Recent studies like SegPoint (He et al. 2024) and Reason3D (Huang et al. 2025) integrate LLMs’ reasoning ability into 3D segmentation, but they are limited to segmentation for single-category objects and have no textual explanations. In contrast, our approach targets multi-object 3D segmentation and provides textual explanations in the output answer, enhancing the model’s understanding of 3D spatial relations and offering a more practical solution for real-world complex scenes.
Large Multimodal Models
Inspired by the remarkable learning ability of Large Language Models (LLMs), recent research has expanded into the visual domain and developed a series of Large Multimodal Models (LMMs) (Alayrac et al. 2022; Zhang, Li, and Bing 2023). The prevalent approach focuses on aligning visual representations with the linguistic embeddings of LLMs. For example, BLIP-2 (Li et al. 2023) and mPLUG-OWL (Ye et al. 2023) encode image features with a visual encoder, integrating them into the LLMs with text embeddings. LLaVA (Liu et al. 2023a) and MiniGPT4 (Zhu et al. 2024b) align image-text features followed by instruction tuning, and they also explore image retrieval for LLMs. Recent studies delve into the integration of multimodal LLMs with vision tasks. For example, VisionLLM (Wang et al. 2023) provides an interface for vision-centric tasks via instruction tuning though it does not exploit LLMs for complex reasoning. VisionLLM-v2 (Wu et al. 2024d) integrates visual perception, understanding, and generation within a unified framework by using a “super link” to connect the multimodal large model with task-specific decoders. DetGPT (Pi et al. 2023) introduces multimodal LLMs into open-vocabulary detectors for instruction-based detection tasks. GPT4RoI (Zhang et al. 2023) introduces spatial boxes as inputs, training on region-text pairings. LISA (Lai et al. 2024) enhances segmentation in multimodal LLMs by introducing a <SEG> token. These existing studies primarily target downstream tasks in the 2D domain. In contrast, our approach extends into the 3D domain to enable multi-object segmentation of 3D point clouds and provides textual explanations to support reasoning in the segmentation process.
ReasonSeg3D Dataset
Most existing reasoning segmentation datasets are not suitable for studying the proposed 3D multi-object reasoning segmentation task. Specifically, existing reasoning segmentation datasets have two critical limitations. First, the existing 2D reasoning segmentation datasets (Lai et al. 2024; Ren et al. 2024; Wang and Ke 2024; Yan et al. 2024a) lack 3D data, making them unsuitable for 3D tasks. Second, the existing 3D reasoning datasets (He et al. 2024; Huang et al. 2025) focus on single-category objects as illustrated in Table 1, and they contain only a few hundred scenes which are far fewer than standard 3D segmentation datasets (Dai et al. 2017; Rozenberszki, Litany, and Dai 2022). Additionally, these 3D reasoning datasets lack textual explanations with spatial information, hindering them from training MLLMs for better 3D spatial relation understanding. We bridge this gap by proposing a data generation pipeline as well as a new 3D reasoning segmentation dataset ReasonSeg3D, with more details to be elaborated in the following subsections.
Dataset Definition
@@
In the proposed ReasonSeg3D, each point cloud $P$ is paired with multiple $\{x_{que},y_{ans},M\}$ triplets, where $x_{que}$ is a question targeting one or more objects in the image, $y_{ans}$ is the textual answer containing the explanation for the reasoning segmentation, and $M$ represents the 3D segmentation masks corresponding to $y_{ans}$. The question $x_{que}$ is designed to require world knowledge and reasoning ability to accurately identify and segment multiple objects. For example, instead of directly asking “Where are the sofa and table?”, ReasonSeg3D formulates the question by “Where would be the most suitable place for reading a book in this layout?”. An answer $y_{ans}$ in ReasonSeg3D is constructed to involve multiple objects, and it also includes explanations with 3D spatial relations. For example, the $y_{ans}$ to the above question is formulated by “The corner with the sofa and a small table next to it can serve as a perfect reading nook, providing comfort and a quiet atmosphere.”.
@@
Figure 2: Overview of our proposed MORE3D: Given an input point cloud, the 3D Encoder first extracts per-point features $F_{p}$ and projects them into sequential features $F_{s}$. The sequential features $F_{s}$, together with the textual input $X_{que}$, are then fed into a multimodal LLM to perform reasoning, obtaining LLM embeddings and then producing textual answers $\hat{Y}_{ans}$ with both detailed explanations and descriptions of 3D spatial relations among multiple objects. Finally, after object-specific embedding extraction, object embeddings for multiple <SEG> tokens $F_{seg}$ and the per-point features $F_{p}$ are passed to the 3D Decoder to produce 3D segmentation masks. Snowflake icon denotes frozen module, and flame icons denote trainable ones.
Dataset Generation Pipeline
Several reasoning segmentation datasets (Lai et al. 2024; Wang and Ke 2024) employ LLaVA (Liu et al. 2023a) for image captioning and text-only GPT-4 to produce question-answer pairs according to generated captions. However, GPT-4 cannot understand the image content well and its generated question-answer pairs often lack crucial 3D spatial relations that are essential for 3D reasoning segmentation in various real-world tasks.
We design a novel data generation pipeline that introduces GPT-4o which possesses superior visual content understanding capabilities. The pipeline allows the generation of more practical questions and answers by incorporating spatial relations among objects in scenes. Specifically, we employ scene image and ground-truth segmentation to enhance GPT-4o’s 3D spatial understanding capabilities. The prompt template is structured in two parts: the first part contains basic requirements for the GPT-4o and the second specifies detailed requirements for generating questions and answers with a focus on describing 3D spatial relations among objects. With this template, GPT-4o autonomously selects objects to form question-answer pairs that reflect the scene’s content and spatial layout. On top of that, we conduct human verification to eliminate errors, such as incorrect object references and spatial‑relation errors, from the generated question–answer pairs. Since GPT-4o demonstrates excellent generation quality, only about 3% of the generated question–answer pairs require manual correction. Samples that maximize diversity in both target objects and spatial‑relation descriptions are then selected, thereby avoiding bias and ensuring high‑quality, varied question–answer pairs.
Dataset Statistics
ReasonSeg3D comprises 1,513 scenes and 20,113 data samples in total, with point clouds sourced from ScanNetv2 (Dai et al. 2017). Following (Dai et al. 2017), we divide ReasonSeg3D into training and validation sets, comprising 1,201 and 312 scenes, respectively. In ReasonSeg3D, each scene has 13.3 questions on average, with 20 different object categories for segmentation. Our dataset focuses on indoor scenes, which are critical for a wide range of real-world embodied applications, such as household robotics and augmented reality, where accurate reasoning and perception in complex, cluttered environments are essential.
Method
Task Definition
@@
Overall Framework
Figure 2 shows the framework of the proposed MORE3D which enables Multi-Object 3D Reasoning segmentation and provides textual Explanations based on the user’s input question. Given an input point cloud $P$, the 3D Encoder first groups and extracts per-point features $F_{p}$ and then projects $F_{p}$ into sequential features $F_{s}$ that can be processed by LLMs. The sequential features $F_{s}$ and input text queries are processed by the LLM, obtaining LLM embeddings and then generating textual answers $\hat{Y}_{ans}$ that include textual explanations $\hat{Y}_{text}$ and multi-object <SEG> tokens $\hat{Y}_{\texttt{<SEG>}}$. The <SEG> tokens $\hat{Y}_{\texttt{<SEG>}}$ indicate the request for 3D segmentation masks, and the corresponding object embeddings $F_{seg}$ are extracted from the LLM embeddings. These object embeddings $F_{seg}$ are further combined with per-point features $F_{p}$ within 3D Decoder by dot product, ensuring that the object-specific information interacts with the overall point cloud features to generate the final 3D segmentation masks $\hat{M}$ and classification predictions.
@@
Different from prior studies on 2D reasoning segmentation (Lai et al. 2024; Ren et al. 2024; Wang and Ke 2024) relying solely on 2D representations, which cannot effectively capture 3D geometric information or spatial relations in complex 3D scenes, MORE3D is tailored for 3D tasks and features a 3D-specific encoder-decoder for capturing complex 3D information. In addition, MORE3D extracts object-specific point cloud embeddings to handle multiple objects in 3D scenes, resolving unique challenges in 3D reasoning segmentation effectively. Further, MORE3D generates textual explanations that incorporate 3D spatial relations among objects, enabling explainable and spatially grounded reasoning of 3D scenes.
Multi-Object 3D Reasoning Segmentation
We design an object-specific embedding extraction approach to achieve multi-object 3D reasoning segmentation. This approach enables precise segmentation and reasoning for multiple objects of different categories in complex point clouds, integrating the LLM with a 3D point cloud decoder to generate segmentation masks and textual explanations.
The LLM takes point cloud features $F_{s}$ and text queries $X_{que}$ as input, generating textual answers $\hat{Y}_{ans}$, which include textual explanations $\hat{Y}_{text}$ and multi-object <SEG> tokens $\hat{Y}_{\texttt{<SEG>}}$. Specifically, each output textual answer consists of a paragraph of textual explanations accompanied by a set of <SEG> tokens, where the number of <SEG> tokens equals the number of 3D objects indicated in the input text query. In each answer, the 3D object’s name is followed by a <SEG> token. For example, the LLM might generate the following answer: “You can use the spacious sofa <SEG> for seating, positioned near a central table <SEG> for drinks and snacks, while additional chairs <SEG> provide extra seating options for guests.” Each <SEG> token indicates a request for a point cloud segmentation mask.
@@
    Method     Venue     cIoU     gIoU
    3D Segmentation Methods
    OpenScene (Peng et al. 2023)     CVPR 23     7.69     9.52
    PLA (Ding et al. 2023)     CVPR 23     10.76     10.27
    RegionPLC (Yang et al. 2024b)     CVPR 24     10.82     11.06
    3D-STMN (Wu et al. 2024c)     AAAI 24     19.37     20.70
    RefMask3D (He and Ding 2024)     MM 24     20.53     22.81
    MDIN (Wu et al. 2024b)     MM 24     23.04     23.49
    3D General Scene Understanding Methods
    3D-LLM (Hong et al. 2023)     NeurIPS 23     21.36     22.49
    3D-VisTA (Zhu et al. 2023)     ICCV 23     23.95     24.77
    3D Reasoning Segmentation Methods
    Reason3D (Huang et al. 2025)     3DV 25     25.12     25.90
    MORE3D (Ours)     -     30.19     32.01
Table 2: Benchmarking on the ReasonSeg3D validation set. Best in bold, second underlined.
Explainability
The LLM-generated textual answers $\hat{Y}_{ans}$ contain explanations of the implicit intention of the user’s questions. In addition, the answer includes the descriptions of 3D spatial relations of multiple objects which provide useful guidance for identifying these objects. Unlike most existing reasoning-based segmentation methods (Wang and Ke 2024; Yan et al. 2024a; He et al. 2024; Huang et al. 2025) that just output statements like "It is <SEG>" without further explanation, our generated answer integrates detailed explanations about 3D spatial relations to enhance segmentation. The ground truth answers $Y_{ans}$ also incorporate these 3D spatial relations and are used to supervise the generated output textual answers, guiding the model to capture spatial information effectively. The textual answer loss is defined by:
@@
where $M$ denotes the ground truth segmentation mask, $\hat{M}$ denotes the predicted segmentation mask, $BCE$ denotes the binary cross-entropy loss and $DICE$ denotes the Dice loss (Milletari, Navab, and Ahmadi 2016). We use the cross-entropy loss to supervise the point cloud classification as follows:
@@
Experiment
Experimental Settings
@@
Following prior studies on reasoning segmentation (Lai et al. 2024; Ren et al. 2024; Wang and Ke 2024), we adopt two evaluation metrics including cumulative IoU (cIoU) and generalized IoU (gIoU). cIoU is computed as the cumulative intersection over cumulative union, while gIoU is the average Intersection-over-Union (IoU) across all samples.